# Qipeng Kuang Ondřej Kuželka Yuanhong Wang Yuyi Wang

### v1

@@ 2.3 Weighted First Order Model Counting
The weighted first-order model counting problem (WFOMC) takes the input consisting of a first-order sentence $\Psi$, a domain size $n\in\mathbb{N}$, and a pair of weighting functions $(w,\overline{w})$ that both map all predicates in $\Psi$ to real weights. Given a set $L$ of true or false ground atoms (i.e., ground literals) whose predicates are in $\Psi$, the weight of $L$ is defined as $W(L,w,\overline{w}):=\prod_{a\in L_{T}}w(\mathsf{pred}\left(a\right))\cdot\prod_{a\in L_{F}}\overline{w}(\mathsf{pred}\left(a\right)),$ where $L_{T}$ (resp. $L_{F}$) denotes the set of true ground (resp. false) atoms in $L$, and $\mathsf{pred}\left(a\right)$ maps an atom $a$ to its corresponding predicate name.
@@
Since these weightings are defined on the predicate level, all groundings of the same predicate get the same weights. For this reason, the notion of WFOMC defined here is also referred to as symmetric WFOMC.
@@
The weighted model counting problem (WMC) is defined similarly but the input formula is a ground propositional sentence.
@@
The WMC of a ground propositional sentence $\Phi$ under weighting functions $(w,\overline{w})$ is
@@
The WMC of a ground propositional sentence $\Phi=R(a,b)\to S(a)$ under $w(R)=2,\overline{w}(R)=w(S)=\overline{w}(S)=1$ is
@@ 2.4 Data Complexity of WFOMC
A sentence, or class of sentences, that enjoys polynomial-time data complexity is said to be domain-liftable.
Domain-liftability of $\text{FO}^{2}$
Building on [dB11, dBMD14], in  [BdBGS15, Appendix C], it was shown that any sentence with at most two logical variables ($\text{FO}^{2}$) is domain-liftable by devising a polynomial-time algorithm for computing $\mathsf{WFOMC}(\Psi,n,w,\overline{w})$ for any $\text{FO}^{2}$ sentence $\Psi$ and any weighting $(w,\overline{w})$.
@@
Let $C_{1}(x),C_{2}(x),\cdots,C_{L}(x)$ be all the 1-types of $\Psi$. When we consider domain-liftability, the sentence $\Psi$ is fixed and therefore $L$ is a constant. We partition domain elements into 1-types by enumerating the configuration $\bm{c}=(c_{1},c_{2},\cdots,c_{L})$, a vector indicating the number of elements partitioned to each 1-type, since in the symmetric WFOMC only the configuration matters. Then we can rewrite $\Psi$ as
$\eta=$ $\bigwedge_{1\leq i<j\leq L}\forall x\in C_{i}\forall y\in C_{j}(\psi(x,y)\land\psi(y,x))$
$\land\bigwedge_{1\leq i\leq L}\forall x\in C_{i}\forall y\in C_{i}\ \psi(x,y).$
Since we know the truth values of the unary and reflexive binary atoms of each 1-type, we may simplify the body of each conjunct by substituting every unary and reflexive binary atom with true or false as appropriate. Denote by $\psi_{i}(x,y)$ the simplification of $\psi(x,y)$ when both x and y belong to the same 1-type $C_{i}$, and by $\psi_{ij}(x,y)$ the simplification of $\psi(x,y)\land\psi(y,x)$ when $x$ and $y$ belong to $C_{i}$ and $C_{j}$ respectively. We then have:
$\eta=$ $\bigwedge_{1\leq i<j\leq L}\forall x\in C_{i}\forall y\in C_{j}\ \psi_{ij}(x,y)$
$\land\bigwedge_{1\leq i\leq L}\forall x\in C_{i}\forall y\in C_{i}\ \psi_{i}(x,y).$
An important property here is that each conjunct is independent of others since no two conjuncts involve the same ground binary atom. Let
@@
$s_{i}$ $=\mathsf{WMC}(\psi_{i}(a,b)\land\psi_{i}(b,a),w,\overline{w}).$
Then we have
@@
Domain-liftability of Extended $\text{C}^{2}$
@@
Definition 2.5 (Extended $\text{C}^{2}$).
An extended $\text{C}^{2}$ sentence is a $\text{C}^{2}$ sentence possibly with cardinality constraints and unary evidence.
Counting quantifiers, cardinality constraints, and unary evidence are useful to encode certain properties of graphs. An example is using cardinality constraints to encode the identity relation $Eq$ as mentioned above. We present another example of using unary evidence to encode a block-structured graph in the following.
@@
A block-structured graph is an undirected graph $G=(V,E)$ where vertices are partitioned into a constant number of blocks, and for every two blocks $V_{1},V_{2}$, either $\forall x\in V_{1}\forall y\in V_{2},(x,y)\in E$ or $\forall x\in V_{1}\forall y\in V_{2},(x,y)\not\in E$. With unary evidence, we can encode a block-structured graph in the following way: we introduce a fresh unary predicate $Block_{i}$ for each block $V_{i}$, and let
@@
$\land\bigwedge_{i=1}^{n}\left(Block_{v(i)}(i)\land\bigwedge_{j\neq v(i)}\lnot Block_{j}(i)\right),$
where $v(i)$ indicates the block that vertex $i$ belongs to, and $Block_{*}(i)$ and $\lnot Block_{*}(i)$ are unary evidence for vertex $i$.
@@
When the input sentence of WFOMC involves cardinality constraints and unary evidence, i.e., the sentence is in extended $\text{C}^{2}$, we need to analyze the data complexity more carefully. In such instances, the $\text{C}^{2}$ sentence in the input remains fixed, while the parameters of the constraints and the number of facts in the unary evidence, which might depend on the domain size, are considered variable inputs.
The presence of counting quantifiers, cardinality constraints, and unary evidence does not change the domain-liftability of $\text{FO}^{2}$ even if the parameters of cardinality constraints and the length of unary evidence depend on the domain size. By the techniques in [Kuz21] and [WPWK24, Appendix A], the WFOMC of an extended $\text{C}^{2}$ sentence can be reduced to the WFOMC of a sentence in $\text{FO}^{2}$ with symbolic weightings, and then solved by the algorithm for $\text{FO}^{2}$. As we will see, all these techniques apply to the methods in this paper. Therefore, in the rest of this paper, we only present the algorithms for $\text{FO}^{2}$, and the results can be extended to extended $\text{C}^{2}$.
@@ 2.5 Axioms
While extended $\text{C}^{2}$ captures numerous graph properties, it falls short in expressing certain important characteristics, such as trees, forests, and strongly connected orientations. In fact, these structures cannot be finitely axiomatized by first-order logic without grounding them out. However, in terms of WFOMC, it is possible to express these properties more succinctly using what we refer to as axioms.
@@ 2.6 Graph Polynomial
Graph polynomials are a class of polynomials that are defined on graphs and have combinatorial meaning.
Tutte Polynomial
The Tutte polynomial [Tut54] is one of the most well-known graph polynomials for undirected graphs. Consider an undirected graph $G=(V,E)$, where $V$ is the set of vertices and $E$ is the set of edges. The Tutte polynomial for the graph $G$ is a polynomial in two variables $x$ and $y$ defined as
@@
Directed Chromatic Polynomial
The idea of the chromatic polynomial of a directed graph appeared in [Sta70] though they stated it in terms of ordered sets. Given a directed graph $D=(V,E)$ where $n=|V|$, let $\chi_{D}(x)$ be the number of ways of coloring each vertex in $D$ by one of the colors $\{1,2,\cdots,x\}$ such that if there is an edge from $u$ to $v$, the color of $u$ is smaller than the color of $v$. We have the following lemma to state that $\chi_{D}(x)$ is a polynomial of $x$.
@@ 3 Weak Connectedness Polynomial
In this section, we introduce our first polynomial, called Weak Connectedness Polynomial (WCP). The WCP is a univariate polynomial defined on a first-order logic sentence and a distinguished binary relation. Similarly to graph polynomials, the WCP can also capture various combinatorial properties of the models of the sentence w.r.t. the binary relation. As we show later in this section, WCP can be computed in polynomial time in the domain size for the $\text{C}^{2}$ fragment with cardinality constraints and unary evidence. This property allows us to attach axioms of non-trivial combinatorial structures and complicated weight constraints to the first-order logic sentence and derive Tutte polynomial for certain graphs.
@@ 3.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and unary evidence, $w,\overline{w}$ be two weighting functions, and $R$ be a distinguished binary relation. The $n$-th WCP $f_{n}(u)$ of $\Psi$ for the relation $R$ is the unique univariate polynomial which satisfies:
@@
By $\bigwedge_{i=2}^{u}\forall x\left(A_{i}(x)\to A_{i-1}(x)\right)$, for each element $e$ in the domain there must be an $i^{\prime}\in[0,u]$ such that $A_{1}(e)=\cdots=A_{i^{\prime}}(e)=\top$ and $A_{i^{\prime}+1}(e)=\cdots A_{u}(e)=\bot$. Meanwhile, elements whose corresponding vertices form a weakly connected component in $G(\mu_{R})$ must have the same $i^{\prime}$ since
$\bigwedge_{i=1}^{u}\forall x\forall y\left(A_{i}(x)\wedge R(x,y)\to A_{i}(y)\right).$
@@
As WCP is defined based on WFOMC, it can be also extended to first-order logic sentences with counting quantifiers, cardinality constraints, and unary evidence. The counting quantifiers and unary evidence can be naturally incorporated into the first-order logic sentences. Besides, cardinality constraints are usually treated as atomic formulas in the previous WFOMC literature [Kuz21, MBS23, vBK21], where the satisfaction relation as well as the model definition are extended accordingly.
Example 3.1.
Consider the sentence encoding $k$-regular graphs:
$\Phi=\forall x\lnot E(x,x)\land\forall x\forall y\left(E(x,y)\to E(y,x)\right)\land\forall x\exists_{=k}yE(x,y).$
By Proposition 3.1, its $n$-th WCP $f_{n}(u;\Psi,\mathds{1},\mathds{1},E)$ is $\sum_{g\in G(n,k)}(u+1)^{cc(g)}$, where $\mathds{1}$ is the constant weighting function that always returns $1$, and $G(n,k)$ is the set of $k$-regular graphs with $n$ vertices. The coefficients $[u^{i}]f_{n}(u-1;\Psi,\mathds{1},\mathds{1},E)$ count the number of $k$-regular graphs with $i$ connected components.
@@ 3.2 Fast Calculation
Note that by definition, evaluating the polynomial at a specific positive integer is a WFOMC query. However, such WFOMC is not readily solvable by existing algorithms (e.g., [BdBGS15, vBK21]) even for the $\text{FO}^{2}$ fragment. In general, the runtime of these algorithms is exponential in the length of the sentence, while the length of the sentence $\Psi_{R,u}$ in the WFOMC query can be as large as the domain size $n$, hence these algorithms would not lead to polynomial runtime in the domain size. In particular, the algorithm described in Section 2.4 is not sufficient in this case either, as there would be $O(2^{n})$ 1-types and, consequently, an exponential time data complexity. To this end, we adapt this algorithm into a dynamic programming (DP) style algorithm that exploits the structure of $\Psi_{R,u}$ to achieve an efficient calculation of the polynomials for $\text{FO}^{2}$ sentences.
Let $\Psi$ and $R$ be the input $\text{FO}^{2}$ sentence and the binary relation that WCP is defined on. Consider the point evaluation of the WCP at $u$. Let $C_{1}(x),C_{2}(x),\cdots,C_{L}(x)$ be the valid 1-types  [^6] We call a 1-type valid if it is satisfiable in some model of the sentence. of $\Psi$. Due to the definition of $\Psi_{R,u}$, the valid 1-types of $\Psi_{R,u}$ must have the form $C_{i}(x)\wedge C^{A}_{i^{\prime}}(x)$, where $C^{A}_{i^{\prime}}(x)$ is defined as
@@
Observation 3.1.
Reformulate the WFOMC of $\Psi_{R,u}$ to the WFOMC of a sentence $\forall x\forall y\ \phi(x,y)$ in prenex normal form with only universal quantifiers (removing existential quantifiers using the Skolemization technique [dBMD14]). Let $\phi_{(i,i^{\prime}),(j,j^{\prime})}(x,y)$ be the simplification of $\phi(x,y)\land\phi(y,x)$ when $x$ and $y$ belong to 1-type $(i,i^{\prime})$ and $(j,j^{\prime})$ respectively. Define
$r_{(i,i^{\prime}),(j,j^{\prime})}=\mathsf{WMC}(\phi_{(i,i^{\prime}),(j,j^{\prime})}(a,b),w,\overline{w}).$
It holds that
@@
This observation provides an important insight into the structure of the new 1-types (as shown in Figure 1): we can imagine that the new 1-types of $\Psi_{R,u}$ form $u+1$ layers, with each layer containing a replica of the original 1-types of $\Psi$. If two elements $a,b$ fall in 1-types in the same layer, $R(a,b)$ can be either true or false. But if $a$ and $b$ are located in 1-types in different layers, $R(a,b)$ as well as $R(b,a)$ must be false due to the definition of $\Psi_{R,u}$ in Equation 3.
@@
Figure 1: The 1-type structure of $\Psi_{R,u}$. Each tiny gray box represents an original 1-type of $\Psi$. The bold black cells refer to the replica of the original 1-types in each layer. The DP algorithm proceeds from left (red cells) to right (blue cell).
@@
is the weight between the layer $\hat{u}$ and the layers before $\hat{u}$. The initial values are $h_{0,\bm{c}}=w_{in}(\bm{c})$ for each $\bm{c}$. The final WFOMC of $\Psi_{R,u}$ can be obtained by:
@@
Input: Sentence $\Psi$, weighting functions $w,\overline{w}$, the relation $R$, domain size $n$
@@
Computing WCP for a given extended $\text{C}^{2}$ sentence, a domain of size $n$, a distinguished binary relation, and weighting functions can be computed in polynomial time in $n$.
@@
As discussed in Section 2.4, any WFOMC problem for extended $\text{C}^{2}$ sentences can be reduced to WFOMC of an $\text{FO}^{2}$ sentence, which can be solved in polynomial time in $n$ by the algorithm presented above. The interpolation step from $n+1$ point evaluations to the polynomial can also be done in polynomial time in $n$, completing the proof. ∎

### v2

@@ 2.3 Weighted First Order Model Counting
The weighted first-order model counting problem (WFOMC) takes the input consisting of a first-order sentence $\Psi$, a domain size $n\in\mathbb{N}$, and a pair of weighting functions $(w,\overline{w})$ that both map all predicates in $\Psi$ to real weights. Given a set $\mathcal{L}$ of ground literals whose predicates are in $\Psi$, the weight of $\mathcal{L}$ is defined as $W(\mathcal{L},w,\overline{w}):=\prod_{l\in\mathcal{L}_{T}}w(\mathsf{pred}\left(l\right))\cdot\prod_{l\in\mathcal{L}_{F}}\overline{w}(\mathsf{pred}\left(l\right)),$ where $\mathcal{L}_{T}$ (resp. $\mathcal{L}_{F}$) denotes the set of positive (resp. negative) ground literals in $\mathcal{L}$, and $\mathsf{pred}\left(l\right)$ maps a literal $l$ to its corresponding predicate name.
@@
Since these weightings are defined on the predicate level, all positive ground literals of the same predicate get the same weights, and so do all negative ground literals of the same predicate. For this reason, the notion of WFOMC defined here is also referred to as symmetric WFOMC.
@@
The weighted model counting problem (WMC) is defined similarly but the input formula is a ground quantifier-free formula.
@@
The WMC of a ground quantifier-free formula $\Phi$ under weighting functions $(w,\overline{w})$ is
@@
The WMC of a ground quantifier-free formula $\Phi=R(a,b)\to S(a)$ under $w(R)=2,\overline{w}(R)=w(S)=\overline{w}(S)=1$ is
@@ 2.4 Data Complexity of WFOMC
A sentence, or class of sentences, that enjoys polynomial-time data complexity, i.e., the complexity is polynomial in the domain size, is said to be domain-liftable.
Domain-liftability of $\text{FO}^{2}$. Building on [dB11, dBMD14], in  [BdBGS15, Appendix C], it was shown that any sentence with at most two logical variables ($\text{FO}^{2}$) is domain-liftable by devising a polynomial-time algorithm for computing $\mathsf{WFOMC}(\Psi,n,w,\overline{w})$ for any $\text{FO}^{2}$ sentence $\Psi$ and any weighting $(w,\overline{w})$.
@@
Let $C_{1}(x),C_{2}(x),\cdots,C_{L}(x)$ be all the 1-types of $\Psi$. When we consider domain-liftability, the sentence $\Psi$ is fixed and therefore $L$ is a constant. We partition domain elements into 1-types by enumerating the configuration $\bm{c}=(c_{1},c_{2},\cdots,c_{L})$, a vector indicating the number of elements partitioned to each 1-type, since in the symmetric WFOMC only the configuration matters. Then we can rewrite $\Psi$ in the form of
$\Psi=\bigwedge_{1\leq i<j\leq L}\forall x\in C_{i}\forall y\in C_{j}(\psi(x,y)\land\psi(y,x))\land\bigwedge_{1\leq i\leq L}\forall x\in C_{i}\forall y\in C_{i}\ \psi(x,y),$
where $\psi(x,y)$ is a quantifier-free formula.
Since we know the truth values of the unary and reflexive binary atoms of each 1-type, we may simplify the body of each conjunct by substituting every unary and reflexive binary literal with true or false as appropriate. Denote by $\psi_{i}(x,y)$ the simplification of $\psi(x,y)$ when both $x$ and $y$ belong to the same 1-type $C_{i}$, and by $\psi_{ij}(x,y)$ the simplification of $\psi(x,y)\land\psi(y,x)$ when $x$ and $y$ belong to $C_{i}$ and $C_{j}$ respectively. We then have:
$\Psi=\bigwedge_{1\leq i<j\leq L}\forall x\in C_{i}\forall y\in C_{j}\ \psi_{ij}(x,y)\land\bigwedge_{1\leq i\leq L}\forall x\in C_{i}\forall y\in C_{i}\ \psi_{i}(x,y).$
An important property here is that each conjunct is independent of others since no two conjuncts involve the same ground binary literal. Let
@@
$s_{i}$ $=\mathsf{WMC}(\psi_{i}(a,b)\land\psi_{i}(b,a),w,\overline{w})$
be the mutual and internal 1-type coefficient of $\Psi$ respectively. Then we have
@@
Domain-liftability of $\text{C}^{2}$. The above algorithm also serves as the basis of a WFOMC algorithm for $\text{C}^{2}$ (the $\text{FO}^{2}$ sentence with counting quantifiers) together with cardinality constraints and unbounded ground unary literals.
Definition 2.8 (Counting quantifiers).
The counting quantifier $\exists_{=k}$ restricts that the number of assignments of the quantified variable satisfying the subsequent formula is exactly $k$. Counting quantifiers $\exists_{\leq k}$ and $\exists_{\geq k}$ are defined similarly.
Counting quantifiers allow us to express concepts such as “each vertex in the undirected graph has exactly $k$ edges”, i.e., $k$-regular graphs, by the sentence $\forall x\lnot E(x,x)\land\forall x\forall y\left(E(x,y)\to E(y,x)\right)\land\forall x\exists_{=k}yE(x,y)$.
Definition 2.9 (Cardinality constraints).
@@
For example, $|Eq|=n$ means that there are exactly $n$ positive ground literals of $Eq$ in a model, and the sentence $(\forall x\ Eq(x,x))\land|Eq|=n$ along with a domain of size $n$ encodes an identity relation $Eq$. Note that we allow cardinality constraints to depend on the domain size (such as in the definition of $Eq$ here where we used $|Eq|=n$).
Finally, we show the usage of ground unary literals defined in Section 2.2. For example, the sentence $\forall x\exists y(R(x,y)\to S(y))\land S(1)\land\lnot S(2)$ contains ground unary literals $S(1)\land\lnot S(2)$ indicating that any interpretation of this sentence should satisfy that $S(1)$ is true and $S(2)$ is false. Another example is encoding a block-structured graph.
@@
A block-structured graph is an undirected graph $G=(V,E)$ where vertices are partitioned into a constant number of blocks, and for every two blocks $V_{i},V_{j}$, either $\forall x\in V_{i}\forall y\in V_{j},(x,y)\in E$ or $\forall x\in V_{i}\forall y\in V_{j},(x,y)\not\in E$. We can encode a block-structured graph in the following way: we introduce a fresh unary predicate $Block_{i}$ for each block $V_{i}$, and let
@@
$\land\bigwedge_{e=1}^{n}\left(Block_{v(e)}(e)\land\bigwedge_{j\neq v(e)}\lnot Block_{j}(e)\right),$
where $v(e)$ indicates the block that vertex $e$ belongs to, and $Block_{*}(e)$ and $\lnot Block_{*}(e)$ are ground unary literals for vertex $e$ indicating whether $e$ is in the specified block or not.
@@
We need to define the data complexity more carefully when the input sentence of WFOMC involves cardinality constraints and ground unary literals. In such instances, the $\text{C}^{2}$ sentence (including the parameters in counting quantifiers) in the input remains fixed, while cardinality constraints and ground unary literals are considered variable inputs encoded in binary.
The presence of counting quantifiers, cardinality constraints, and ground unary literals does not affect the domain-liftability of $\text{FO}^{2}$. Using the techniques in [Kuz21] and [WPWK24, Appendix A], WFOMC of $\text{C}^{2}$ with cardinality constraints and unbounded ground unary literals can be reduced to WFOMC of $\text{FO}^{2}$ with symbolic weights [^6] The weighting functions $w,\overline{w}$ map predicates to symbolic variables. For more details, see [Kuz21, Section $3.1$]., and then solved by the algorithm for $\text{FO}^{2}$ in a straightforward manner.
Remark 2.12.
All the techniques mentioned above apply to the methods in this paper. Therefore, in the rest of this paper we only present the algorithms for $\text{FO}^{2}$ and state the results in terms of $\text{C}^{2}$ with cardinality constraints for simplicity.
@@ 2.5 Axioms
While $\text{C}^{2}$ captures numerous graph properties, it falls short in expressing certain important characteristics, such as trees, forests, and strongly connected orientations. In fact, these structures cannot be finitely axiomatized by first-order logic.
@@ 2.6 Graph Polynomial
Graph polynomials are a class of polynomials defined on graphs that carry combinatorial meaning.
Tutte Polynomial. The Tutte polynomial [Tut54] is one of the most well-known graph polynomials for undirected graphs. Consider an undirected graph $G=(V,E)$, where $V$ is the set of vertices and $E$ is the set of edges. The Tutte polynomial for the graph $G$ is a polynomial in two variables $x$ and $y$ defined as
@@
Directed Chromatic Polynomial. The idea of the chromatic polynomial of a directed graph appeared in [Sta70], although it was originally stated in terms of ordered sets. Given a directed graph $D=(V,E)$ where $n=|V|$, let $\chi_{D}(x)$ be the number of ways of coloring each vertex in $D$ by one of the colors $\{1,2,\cdots,x\}$ such that if there is an edge from $u$ to $v$, the color of $u$ is smaller than the color of $v$. We have the following lemma to state that $\chi_{D}(x)$ is a polynomial of $x$.
@@ 3 Weak Connectedness Polynomial
In this section, we introduce our first polynomial, called Weak Connectedness Polynomial (WCP). The WCP is a univariate polynomial defined on a first-order logic sentence and a distinguished binary relation. Similarly to graph polynomials, the WCP can also capture various combinatorial properties of the models of the sentence w.r.t. the binary relation. As we show later in this section, WCP can be computed in polynomial time in the domain size for sentences from the fragment $\text{C}^{2}$ with cardinality constraints. This property allows us to attach axioms of non-trivial combinatorial structures to the first-order logic sentence as well as to derive the Tutte polynomial for certain graphs.
@@ 3.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and ground unary literals, $w,\overline{w}$ be two weighting functions, and $R$ be a distinguished binary relation. The $n$-th WCP $f_{n}(u)$ of $\Psi$ for the relation $R$ is the unique univariate polynomial which satisfies:
@@
By $\bigwedge_{i=2}^{u}\forall x\left(A_{i}(x)\to A_{i-1}(x)\right)$, for each element $e$ in the domain there must be an $i^{\prime}\in[0,u]$ such that $A_{1}(e)=\cdots=A_{i^{\prime}}(e)=\top$ and $A_{i^{\prime}+1}(e)=\cdots=A_{u}(e)=\bot$. Meanwhile, elements whose corresponding vertices form a weakly connected component in $G(\mu_{R})$ must have the same $i^{\prime}$ due to the first row of Equation (3).
@@ 3.2 Fast Calculation
Note that by definition, evaluating the polynomial at a specific positive integer is a WFOMC query. However, such WFOMC is not readily solvable by existing algorithms (e.g., [BdBGS15, vBK21]) even for the $\text{FO}^{2}$ fragment. In general, the runtime of these algorithms is exponential in the length of the sentence, while the length of the sentence $\Psi_{R,u}$ in the WFOMC query can be as large as the domain size $n$, hence these algorithms would not lead to polynomial runtime in the domain size. In particular, the algorithm described in Section 2.4 is not sufficient in this case either, as there would be $O(2^{n})$ 1-types and, consequently, an exponential time data complexity. To this end, we adapt this algorithm into a dynamic programming (DP) style algorithm that exploits the structure of $\Psi_{R,u}$ to achieve an efficient calculation of the polynomials for $\text{FO}^{2}$ sentences, which then applies to $\text{C}^{2}$ sentences with cardinality constraints by techniques mentioned in Section 2.4.
Let $\Psi$ and $R$ be the input $\text{FO}^{2}$ sentence and the binary relation that WCP is defined on. Consider the point evaluation of the WCP at $u$. Let $C_{1}(x),C_{2}(x),\cdots,\linebreak C_{L}(x)$ be the valid 1-types  [^7] We call a 1-type valid if it is satisfiable in some model of the sentence. of $\Psi$. Due to the definition of $\Psi_{R,u}$, the valid 1-types of $\Psi_{R,u}$ must have the form $C_{i}(x)\wedge C^{A}_{i^{\prime}}(x)$, where $C^{A}_{i^{\prime}}(x)$ is defined as
@@
Observation 3.4.
Let $r_{(i,i^{\prime}),(j,j^{\prime})}$ be the mutual 1-type coefficient (defined in Section 2.4) of $\Psi_{R,u}$. It holds that
@@
This observation provides an important insight into the structure of the new 1-types (as shown in Figure 1): we can imagine that the new 1-types of $\Psi_{R,u}$ form $u+1$ layers, with each layer containing a replica of the original 1-types of $\Psi$. If two elements $a,b$ fall in 1-types in the same layer, $R(a,b)$ can be either true or false. But if $a$ and $b$ are located in 1-types in different layers, $R(a,b)$ as well as $R(b,a)$ must be false due to the definition of $\Psi_{R,u}$ in Equation (3).
@@
Figure 1: The 1-type structure of $\Psi_{R,u}$. Each tiny gray box represents an original 1-type of $\Psi$. The bold gray cells (each containing $4\times 4$ boxes) refer to the replica of the original 1-types in each layer. The DP algorithm proceeds from left (cells $0,\cdots,3$) to right (cell $4$).
@@
is the weight between the layer $\hat{u}$ and the layers before $\hat{u}$. The initial values are $h_{0,\bm{c}}=w_{in}(\bm{c})$ for each $\bm{c}$. Then the final WFOMC of $\Psi_{R,u}$ can be obtained by:
@@
Input: A sentence $\Psi$, weighting functions $w,\overline{w}$, the relation $R$, and an integer $n$
@@
Computing WCP for a given $\text{C}^{2}$ sentence with cardinality constraints, a domain of size $n$, a distinguished binary relation, and weighting functions can be completed in polynomial time in $n$.
@@
As discussed in Section 2.4, any WFOMC problem for $\text{C}^{2}$ sentences with cardinality constraints can be reduced to WFOMC of an $\text{FO}^{2}$ sentence, which can be solved in polynomial time in $n$ by the algorithm presented above. The interpolation step from $n+1$ point evaluations to the polynomial can also be done in polynomial time in $n$, completing the proof. ∎